# Change Point Detection in Dynamic Graphs with Decoder-only Latent Space Model

### v3

@@ 5.3 Real Data Experiments
In this section, we apply the proposed method to two real data, and we align the detected change points with significant events for interpretation. In practice, the number and location of change points for real data are typically unknown, so there is no widely accepted ground truth for either the change points or their corresponding events in this unsupervised learning problem. Besides validating the detected change points with significant events as in the literature, we attempt a heuristic approach to compare the results across different detection methods as a supplementary evaluation. Specifically, we fit Degree Corrected Stochastic Block Models (DCSBM) (Karrer & Newman, 2011; Zhao et al., 2012) to the networks between consecutive detected change points, and we evaluate the log-likelihood of out-of-sample networks that were excluded during model fitting. We choose DCSBM, a generalization of Stochastic Block Model (SBM), because SBM is a well known approximation of graphons, which are among the most general network model in the literature (Airoldi et al., 2013; Olhede & Wolfe, 2014; Gao et al., 2015). Moreover, DCSBM does not favor either the proposed or competitor methods in terms of fitting the model. Intuitively, a higher log-likelihood suggests that the detected change points segment the time series in a way that better capture the unchanged patterns within each interval. Additional details on this evaluation procedure are provided in Appendix 7.6.
@@ 5.3.2 Enron Email Data
Based on the results in Figure 12, the primary discrepancies between the proposed and competitor methods are observed at the endpoint of the time span. In particular, four competitor methods have detected change points in February 2002, corresponding to events after Enron filed for bankruptcy in December 2001. While the magnitude exhibits two small spikes on the right in Figure 12, they do not exceed the threshold in red to be declared as change points by our method. The scale of the changes toward the communication patterns among employees may be smaller for the events in February 2002, comparing to the scale of changes toward communication patterns during bankruptcy.
This discrepancy suggests a potential extension to the proposed framework. We could specify the priors for the graph representations as $\bm{z}^{t}\sim\mathcal{N}(\bm{\mu}^{t},\bm{\Sigma}^{t})$ and incorporate the covariance $\bm{\Sigma}^{t}$ into the regularization term of the objective function (2) to adaptively detect the changes at different scales. Additionally, the ADMM procedure would require an update for $\bm{\Sigma}^{t}$, and the quadratic form for the localization method would need to be modified as $(\bm{z}^{t}-\bm{z}^{t-1})^{\top}(\bm{\Sigma}^{t}+\bm{\Sigma}^{t-1})^{-1}(\bm{z}^{t}-\bm{z}^{t-1})\sim\chi_{d}^{2}$ when $\bm{\mu}^{t}-\bm{\mu}^{t-1}=\bm{0}$. We consider this refinement as a direction for future development.

### v4

@@ 5.3 Real Data Experiments
In this section, we apply the proposed method to two real data, and we align the detected change points with significant events for interpretation. In practice, the number and location of change points for real data are typically unknown, so there is no widely accepted ground truth for either the change points or their corresponding events in this unsupervised learning problem. Besides validating the detected change points with significant events as in the literature, we attempt a heuristic approach to compare the results across different detection methods as a supplementary evaluation. Specifically, we fit Degree-Corrected Stochastic Block Models (DCSBM) (Karrer & Newman, 2011; Zhao et al., 2012) to the networks between consecutive detected change points, and we evaluate the log-likelihood of out-of-sample networks that were excluded during model fitting. We choose DCSBM, a generalization of Stochastic Block Model (SBM), because SBM is a well known approximation of graphons, which are among the most general network model in the literature (Airoldi et al., 2013; Olhede & Wolfe, 2014; Gao et al., 2015). Moreover, DCSBM does not favor either the proposed or competitor methods in terms of fitting the model. Intuitively, a higher log-likelihood suggests that the detected change points segment the time series in a way that better capture the unchanged patterns within each interval. Additional details on this evaluation procedure are provided in Appendix 7.6.
@@ 5.3.2 Enron Email Data
Based on the results in Figure 12, the primary discrepancies between the proposed and competitor methods are observed at the endpoint of the time span. In particular, four competitor methods have detected change points in February 2002, corresponding to events after Enron filed for bankruptcy in December 2001. While the magnitude exhibits two small spikes on the right in Figure 12, they do not exceed the threshold in red to be declared as change points by our method. The scale of changes in communication patterns among employees is smaller for the events in February 2002, comparing to the scale of changes in communication patterns during bankruptcy.
Besides lowering the threshold to include these change points, the discrepancy suggests a potential extension to improve the proposed framework. Perhaps we could specify the priors for the graph-level representations as $\bm{z}^{t}\sim\mathcal{N}(\bm{\mu}^{t},\bm{\Sigma}^{t})$ and incorporate both the mean $\bm{\mu}^{t}$ and covariance $\bm{\Sigma}^{t}$ into the penalty term of the objective function. For example, to adaptively adjust the changes at different scales, the inverse of the covariance $\bm{\Sigma}^{t}\in\mathbb{R}^{d\times d}$ may serve as a scaling factor for the difference $\bm{\mu}^{t+1}-\bm{\mu}^{t}\in\mathbb{R}^{d}$, thereby enhancing the magnitudes. By learning the covariances of the graph-level representations and monitoring their consecutive shifts over time, the latent space regularized in this way could potentially capture more subtle variations in the data space. Furthermore, the ADMM procedure would require an update for $\bm{\Sigma}^{t}$, and the quadratic form for the localization method would need to be modified as $(\bm{z}^{t}-\bm{z}^{t-1})^{\top}(\bm{\Sigma}^{t}+\bm{\Sigma}^{t-1})^{-1}(\bm{z}^{t}-\bm{z}^{t-1})\sim\chi_{d}^{2}$ where the inverse of the covariance is also used for standardization. As there are different ways to incorporate the covariances into the regularization term and learning the positive semi-definite covariance via neural network is challenging, we consider this improvement as a potential direction for future development.
@@ 6 Discussion
Code Availability
The codes are available at https://github.com/allenkei/CPD_generative.
Acknowledgments
We thank Mark Handcock and Ying Nian Wu for helpful comments on this work. Also, we are extremely grateful for the constructive comments provided by the anonymous reviewers and Action Editors.